OpenStudy (anonymous):

If y varies directly as x, and y = 25 as x = 5, find y for the x-value 7. Answer 20 25 35 30

jimthompson5910 (jim_thompson5910):

y varies directly as x means that y = k*x

jimthompson5910 (jim_thompson5910):

it tells us that when x = 5, y = 25 so y = k*x y = k*5 ... plug in x =5 25 = k*5 ... plug in y = 25, now solve for k k = ???

jimthompson5910 (jim_thompson5910):

tell me what you get for k

OpenStudy (anonymous):

5...?

jimthompson5910 (jim_thompson5910):

good, k = 5

jimthompson5910 (jim_thompson5910):

so y = k*x turns into y = 5x

jimthompson5910 (jim_thompson5910):

now we want to know the value of y when x = 7, so y = 5x y = 5*7 ... plug in x = 7 y = ??

OpenStudy (anonymous):

35. Thanks I get it now!

jimthompson5910 (jim_thompson5910):

correct, yw